OpenStudy (anonymous):

Simplify 3[32 + 6(3 - 5)] =

OpenStudy (anonymous):

\[3(32+6(3-5)) = 3(32+6(-2)) = 3(32-12) = 3(20) = 60\]

OpenStudy (anonymous):

good job and thanks a lot